OpenStudy (anonymous):

or the following functions, f and g, determine (a) f(g(x)) (b) g(f(x)) If f(x) = 9 x^{3} - 2 x + 11 and g(x) = \sqrt[3]{x}, then (a) f(g(x)) = ; (b) g(f(x)) = .

OpenStudy (anonymous):

\[f(x) = 9 x^{3} - 2 x + 11 , g(x) = \sqrt[3]{x}\]

OpenStudy (anonymous):

\[f(g(x))=f(\sqrt[3]{x})\] is a start

OpenStudy (anonymous):

then \[f(g(x))=9\times \sqrt[3]{x}^3-2\sqrt[3]{x}+11\] \[f(g(x))=9x-2\sqrt[3]{x}+11\]